OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 2 sin^2x = sin x

OpenStudy (anonymous):

x can be either 30 degrees or 150 degrees

OpenStudy (anonymous):

you mean the unit circle right? I always have trouble with that

OpenStudy (anonymous):

want me to show you how?

OpenStudy (anonymous):

I'd love that! By the way this is a multiple choice question...so the answer looks like pi/2 , 3pi/2...like that...

OpenStudy (anonymous):

Does that make it easier?

OpenStudy (anonymous):

Oh and this is also the first time I've ever asked a question on this website so I'm really thankful for the help! ^-^

OpenStudy (zzr0ck3r):

let a=sin(x) you have 2a^2-a=0 a(2a-1)=0 so a=0 or a=1/2 so sin(x) = 0 or sin(x) = 1/2 when does this happen? well sin(x) = 0 when x is 0,pi, and 2pi but we don't include 2pi sin(x) = 1/2 when x is pi/6 and 5pi/6 so x=0,pi,pi/6,5pi/6

OpenStudy (zzr0ck3r):

understand?

OpenStudy (anonymous):

hmm I think so...I guess Death left...he was gonna help me with the unit circle. Can we try one more?

OpenStudy (zzr0ck3r):

the unit circle is just memorization. You just need to know the basic angles. and if you know sin(x) and cos(x) for 0,pi/3,pi/4,pi,6,and pi/2 you can get the rest real easy

OpenStudy (zzr0ck3r):

http://www.youtube.com/watch?v=1-hrT1Ys390 this should help

OpenStudy (anonymous):

7 sin^2x - 14 sin x + 2 = -5 ...but what if it isn't equal to 0?

OpenStudy (zzr0ck3r):

what if x^2+3x=4?

OpenStudy (zzr0ck3r):

you set it to zero right?

OpenStudy (zzr0ck3r):

x^2+3x-4=0

OpenStudy (zzr0ck3r):

so again let a=sin(x) you have 7a^2-14a+7=0

OpenStudy (zzr0ck3r):

do you understand what I did?

OpenStudy (anonymous):

you replaced the sin with a to make it less confusing right?

OpenStudy (anonymous):

ah! didn't mean tot write that you cancelled the -5 by adding it to both sides which gave you the 7

OpenStudy (anonymous):

right?